Chapter 02, Problem 61 Chalkboard Video A cement block accidentally falls from rest from the ledge of a 82.6-m-high building. When the block is 11.3 m above the ground, a man, 1.90 m tall, looks up and notices that the block is directly above him. How much time, at most, does the man have to get out of the way? Number the tolerance is +/-2% Units
Answer #1

velocity at the end of 82.6 - 11.3 m

v = sqrt (2*9.8*71.3)

v = - 37.38 m/s

Now,

11.3 - 1.90 = -37.38t - 4.9t2

9.4 = -37.38t - 4.9t2

4.9t2 + 37.38t - 9.4 = 0

t = 0.2436 seconds